# Distributed Differentially Private Data Analytics via Secure Sketching

### v1

@@ Appendix C Utility Proofs
$\leq(1+\alpha_{S})^{2}\left(\frac{1}{s}\|\sum_{i\in[s]}\mathbf{S}_{i}((\mathbf{A}+\mathbf{G}_{i})\mathbf{x}_{\text{OPT}}-(\mathbf{b}+\mathbf{g}_{i}))\|^{2}+\lambda\|\mathbf{x}_{\text{OPT}}\|^{2}\right)$
@@
(4) $\cdot\left(\frac{1}{s}\|\sum_{i\in[s]}S_{i}((\mathbf{A}+\mathbf{G}_{i})\mathbf{x}_{\text{OPT}}-(\mathbf{b}+\mathbf{g}_{i}))\|^{2}+\lambda\cdot\|\mathbf{x}_{\text{OPT}}\|^{2}\right)$
@@
$\leq$ $(1+\alpha_{S})^{2}\|\mathbf{A}\mathbf{x}_{\text{OPT}}-\mathbf{b}\|^{2}+\left(1+\frac{1}{\alpha_{S}}\right)\eta\sigma^{2}nd\log\frac{1}{\beta}(\|\mathbf{x}_{\text{OPT}}\|^{2}+1)$
$+\lambda\|\mathbf{x}_{\text{OPT}}\|^{2}$
@@
$\leq$ $\left((1+\alpha_{S})^{2}+\left(1+\frac{1}{\alpha_{S}}\right)\frac{\eta\sigma^{2}nd\log\frac{1}{\beta}}{\lambda}\right)\left(\|\mathbf{A}\mathbf{x}_{\text{OPT}}-\mathbf{b}\|^{2}+\lambda\cdot\|\mathbf{x}_{\text{OPT}}\|^{2}\right)$
@@
(5) $\leq$ $\left((1+\alpha_{S})^{4}+\frac{32}{\alpha_{S}}\left(\frac{\eta\sigma^{2}nd\log\frac{1}{\beta}}{\lambda}+\frac{2}{\alpha_{S}}\left(\frac{\eta\sigma^{2}nd\log\frac{1}{\beta}}{\lambda}\right)^{2}\right)\right)$
$\left(\|\mathbf{A}\mathbf{x}_{\text{OPT}}-\mathbf{b}\|^{2}+\lambda\|\mathbf{x}_{\text{OPT}}\|^{2}\right)$
@@
By our choice of $\sigma^{2}$ from Theorem 3.4 and using the sketching matrix of [15] (see Table 5), we have $\sigma^{2}\cdot n\in O(d^{2}\log^{4}d\cdot\alpha_{S}^{-4}\cdot\log^{4}\frac{1}{\beta}\varepsilon^{-2}\cdot\log\frac{1}{\delta})$. Thus, we have
$\leq$ $(\|\mathbf{A}\mathbf{x}_{\text{OPT}}-\mathbf{b}\|^{2}+\lambda\|\mathbf{x}_{\text{OPT}}\|^{2})\cdot\Bigg{(}1+15\alpha_{S}$
$+\tilde{O}\left(\frac{d^{3}\alpha_{S}^{-5}\log^{5}\frac{1}{\beta}\cdot\varepsilon^{-2}\log\frac{1}{\delta}}{\lambda}+\frac{d^{6}\alpha_{S}^{-10}\log^{10}\frac{1}{\beta}\cdot\varepsilon^{-4}\log^{2}\frac{1}{\delta}}{\lambda^{2}}\right)\Bigg{)}$
$+\tilde{O}\left(\frac{d^{3}\alpha_{S}^{-5}\log^{5}\frac{1}{\beta}\cdot\varepsilon^{-2}\log\frac{1}{\delta}}{\lambda}+\frac{d^{6}\alpha_{S}^{-10}\log^{10}\frac{1}{\beta}\cdot\varepsilon^{-4}\log^{2}\frac{1}{\delta}}{\lambda^{2}}\right)$
As a final remark, if $\lambda$ is sufficiently large, one should choose $\alpha_{S}$ to be as small as possible, minimizing the tradeoff between $\alpha_{S}$ and $\frac{\alpha_{S}^{-5}}{\lambda}$. If $\lambda$ is not sufficiently large, one should choose $\alpha_{S}=1/3$ such that $\mathbf{S}$ is an oblivious subspace embedding, but the target dimension does not have a prohibitively large dependency on the sketch distortion.
Appendix D Experimental Evaluation
In this section we provide the results of more combinations of parameters for our utility and running time experiments. We also provide more in depth descriptions of the real-world datasets we used.
D.1. Running Time of LTM
Table 6 provides running times per server and the total communication load in our model for $s=1$ using $3$ servers, and Table 7 shows those when varying $s$.
@@ D.1. Running Time of LTM
D.2. Ridge Regression
On Figure 4, we provide plots for the same setting of parameters as on Figure 3 ($d=10$ and $\mu^{2}=n$). Though, we also vary $\lambda\in\{1,10,100\}$ here. The experiments show that varying $\lambda$ does not change the asymptotic behavior of any of the mechanisms we investigated, in the sense that a variance proportional to $n^{-0.6}$ or lower eventually had decreasing error, while a variance proportional to $n^{0.5}$ produced increasing errors. The increase of $\lambda$ does however produce a significantly lower error in all settings (note the $y$-axes on the figure). Our theoretical results likewise predict this behaviour, as while the error bounds of Theorem 4.4 improve with increasing $\lambda$, they do not affect the dependency on $n$. Thus, we view the experiments as a confirmation that the theoretical bounds, while potentially improvable, express the correct asymptotic relationship between parameters and approximation bounds.
@@ D.2. Ridge Regression
[…]
D.3. Low-Rank Approximation
Figure 5 provides the error for low-rank approximation, where both $\varepsilon$ and $k$ are varied.
Figure 5. Plots depicting the asymptotic behavior of error $\psi$ for $\varepsilon\in\{0.01,0.05\}$ (top, bottom) and $k\in\{5,10\}$ (left, right), with $d=50$. The grey line depicts the error of the local mechanism and the orange one depicts our approach. The other lines resemble different values of $p$. The standard deviations are depicted by the vertical black lines.
[…]
D.4. Real-World Datasets
In addition to synthetic datasets, we also evaluated our mechanism for ridge regression on $4$ datasets from the UC Irvine Machine Learning Repository. The following table provides their number of entries $n$ and dimensionality $d$. A more thorough description of the datasets is found below the table.
Dataset $n$ $d$
Power [34] $2049280$ $6$
Elevation [39] $434874$ $2$
Ethylene [29] $4178504$ $18$
Songs [11] $515345$ $89$
  • The first dataset consists of electric power consumption measurements in one household [34] and the feature we try to predict is sub_metering_3. We ignore the date and time features and the data points that had missing values. This leaves us with $6$ features (plus the one we are predicting) and $2049280$ data points.
  • The Elevation dataset [39] consists of $434874$ open street map elevation measurements from North Jutland, Denmark. We predict the elevation from the longitude and latitude features.
  • The Ethylene dataset [29] contains recordings of sensors exposed to a mixture of gas. We trained on the part where the sensors were exposed to a mixtures of Ethylene and CO in air. The feature we are predicting is the last one TGS2620, which leaves us with $d=18$ and $n=4178504$.
  • The Songs dataset [11], consists of $89$ audio features that are meant to predict the release year of a song.
Appendix E Distributed Computation of Linear Sketches
The linear transformation we consider is motivated by efficient secure distributed cryptographic tools from secure multi-party computation; in particular, we can use linear secret-sharing. We require this secret sharing scheme to be information-theoretically secure. We use an additive secret sharing scheme, such that clients $C_{1},\dots,C_{n}$ generate $k$ shares of their noisy inputs $x^{1},\dots,x^{n}$, where $k$ is the number of servers that compute the linear transformation. We use $[x^{i}]$ to denote a secret sharing of some client $C_{i}$’s noisy input $x^{i}$ for some $i\in[n]$. For some field $\mathbb{F}$ of size $p$ and some prime number $p$, $[x^{i}]$ consists of shares $x^{i}_{1},\dots,x^{i}_{k}\in\mathbb{F}$ such that $\sum_{j=1}^{k}x_{j}^{i}=x^{i}$. To split a secret into $k$ shares, a client can sample $k-1$ random field elements $x^{i}_{1},\dots,x^{i}_{k-1}$ and compute the last secret share as $x^{i}-\sum_{j=1}^{k-1}x_{j}^{i}$. Therefore, it also seems intuitive that an adversary that sees all but one of the shares knows nothing about the input. For every $j\in[k]$, server $S_{j}$ receives $\{x_{j}^{i}\}_{i\in n}$ from the respective parties. We define these shares such that the security of our distributed protocol does not rely on any computational assumption.
Since the linear transformation is public, servers can apply the transformation locally to their secret shares without any need to communicate with each other. Each server then reveals their shares of the resulting linear transformation, such that the linear sketch can be revealed in the clear. For more details on MPC based on secret sharing, see e.g., [Cramer_Damgård_Nielsen_2015].

### v2

@@ 5. Numerical Linear Algebra in LTM
$\leq(1+\alpha_{S})^{2}$
$\left(\frac{1}{s}\|\sum_{i\in[s]}\mathbf{S}_{i}((\mathbf{A}+\mathbf{G}_{i})\mathbf{x}_{\text{OPT}}-(\mathbf{b}+\mathbf{g}_{i}))\|^{2}+\lambda\|\mathbf{x}_{\text{OPT}}\|^{2}\right)$
@@
(9) $\cdot\Bigg{(}\frac{1}{s}\|\sum_{i\in[s]}S_{i}((\mathbf{A}+\mathbf{G}_{i})\mathbf{x}_{\text{OPT}}-(\mathbf{b}+\mathbf{g}_{i}))\|^{2}+\lambda\cdot\|\mathbf{x}_{\text{OPT}}\|^{2}\Bigg{)}$
@@
$\leq$ $(1+\alpha_{S})^{2}\|\mathbf{A}\mathbf{x}_{\text{OPT}}-\mathbf{b}\|^{2}$
$+\left(1+\frac{1}{\alpha_{S}}\right)\eta\sigma^{2}nd\log\frac{1}{\beta}(\|\mathbf{x}_{\text{OPT}}\|^{2}+1)+\lambda\|\mathbf{x}_{\text{OPT}}\|^{2}$
@@
$\leq$ $\left((1+\alpha_{S})^{2}+\left(1+\frac{1}{\alpha_{S}}\right)\frac{\eta\sigma^{2}nd\log\frac{1}{\beta}}{\lambda}\right)$
$\left(\|\mathbf{A}\mathbf{x}_{\text{OPT}}-\mathbf{b}\|^{2}+\lambda\cdot\|\mathbf{x}_{\text{OPT}}\|^{2}\right)$
@@
(10) $\leq$ $\Bigg{(}\frac{32}{\alpha_{S}}\left(\frac{\eta\sigma^{2}nd\log\frac{1}{\beta}}{\lambda}+\frac{2}{\alpha_{S}}\left(\frac{\eta\sigma^{2}nd\log\frac{1}{\beta}}{\lambda}\right)^{2}\right)$
$+(1+\alpha_{S})^{4}\Bigg{)}\left(\|\mathbf{A}\mathbf{x}_{\text{OPT}}-\mathbf{b}\|^{2}+\lambda\|\mathbf{x}_{\text{OPT}}\|^{2}\right)$
@@
By our choice of $\sigma^{2}$ from Theorem 4.7 and using the sketching matrix of (Cohen, 2016) (see Table 2), we have $\sigma^{2}\cdot n\in O(d^{2}\log^{4}d\cdot\alpha_{S}^{-4}\cdot\log^{4}\frac{1}{\beta}\varepsilon^{-2}\cdot\log\frac{1}{\delta})$. Thus, we have
$\|\mathbf{A}\mathbf{x}^{\prime}-\mathbf{b}\|^{2}+\lambda\|x^{\prime}\|^{2}\leq(\|\mathbf{A}\mathbf{x}_{\text{OPT}}-\mathbf{b}\|^{2}+\lambda\|\mathbf{x}_{\text{OPT}}\|^{2})\cdot\Bigg{(}1+15\alpha_{S}$
$+\tilde{O}\Bigg{(}\frac{d^{3}\alpha_{S}^{-5}\log^{5}\frac{1}{\beta}\cdot\varepsilon^{-2}\log\frac{1}{\delta}}{\lambda}+\frac{d^{6}\alpha_{S}^{-10}\log^{10}\frac{1}{\beta}\cdot\varepsilon^{-4}\log^{2}\frac{1}{\delta}}{\lambda^{2}}\Bigg{)}\Bigg{)}$
$+\tilde{O}\Bigg{(}\frac{d^{3}\alpha_{S}^{-5}\log^{5}\frac{1}{\beta}\cdot\varepsilon^{-2}\log\frac{1}{\delta}}{\lambda}+\frac{d^{6}\alpha_{S}^{-10}\log^{10}\frac{1}{\beta}\cdot\varepsilon^{-4}\log^{2}\frac{1}{\delta}}{\lambda^{2}}\Bigg{)}$
Note that we can get a constant multiplicative error as long as the regularization factor $\lambda$ depends on a sufficiently large polynomial in $d$, while being independent of $n$. Generally such a relationship still has good generalization properties when training a regression model. If $\lambda$ is sufficiently large, one should choose $\alpha_{S}$ to be as small as possible, minimizing the tradeoff between $\alpha_{S}$ and $\frac{\alpha_{S}^{-5}}{\lambda}$. If $\lambda$ is not sufficiently large, one should choose $\alpha_{S}=1/3$ such that $\mathbf{S}$ is an oblivious subspace embedding, but the target dimension does not have a prohibitively large dependency on the sketch distortion.
6. Experimental Evaluation
Section 6.1 provides the running time experiments and in Section 6.2 we discuss our utility investigation both for ridge regression and low-rank approximation.
6.1. Running Time of LTM
The only overhead of the LTM over the local model is the execution of the linear transform in MPC, and comes from distributing $\mathbf{A}+\mathbf{G}_{i}\in{\mathbb{R}}^{n\times d}$ among the servers for $i\in[s]$ and then performing matrix multiplications $\mathbf{S}_{i}(\mathbf{A}+\mathbf{G}_{i})$ in MPC, where $\mathbf{S}_{1},\dots,\mathbf{S}_{s}$ is a decomposition of some $\mathbf{S}\in{\mathbb{R}}^{m\times n}$ chosen according to (Nelson and Nguyên, 2013; Cohen, 2016) with sparsity $s$. We implemented our approach in MP-SPDZ (Keller, 2020), a popular and easy to use MPC framework, for multiple combinations of $n$ and $s$, using $S$ servers from Amazon Web Services t3.large instances. We use additive secret sharing over the ring of integers modulo $2^{64}$, which results in no communication between servers, and semi-honest security against an adversary that corrupts all but one servers.
We fix the dimensionality $d=10$ and $m=50$, and vary the number of clients $n\in\{100000,250000,500000,750000,1000000\}$, the sparsity $s\in\{1,10,20,30,40,50\}$ of the sketch (which dictates how many linear transformations we need to apply) and the number of servers $S\in\{2,3\}$. The data matrix $\mathbf{A}$ is generated at random such that all entries are smaller than $2^{32}$ and the decomposition $\mathbf{S}_{1},\dots,\mathbf{S}_{s}$ of some sketch $\mathbf{S}$ is generated according to our mechanism. Every parameter setting is then evaluated by running the protocol $10$ times and averaging over the running times. Table 3 shows the runtime for each server and the total communication load for one particular combination of parameters, which includes three servers: Even with one million clients, the computation on each servers lasts less than $2$ seconds. We conclude that the use of MPC in the LTM does not hinder the computational efficiency of our proposed mechanisms, showing that they can easily be used in practice. Table 4 reports communication and time required when running the protocol with two servers, and Table 5 reports time and communication requirements when varying $s$.
@@ 6.1. Running Time of LTM
6.2. Interpolation between Local and Central DP
Figure 2. Plots depicting the asymptotic behavior of error $\psi$ for $\varepsilon\in\{0.1,0.5\}$ (top, bottom) and $k\in\{5,10\}$ (left, right), with $d=50$. The gray line depicts the error of the local mechanism, and the orange and lime lines depict our approach using Gaussian and Laplacian noise respectively. The other lines resemble different values of $p$ when using Gaussian noise. The standard deviations are depicted by the vertical black lines and the $x$-axis is logarithmic in the number of clients $n$.
Table 6. Parameters of real-world datasets.
Dataset $n$ $d$ $k$
Power (Hebrail and Berard, 2012) $2049280$ $6$ $3$
Elevation (Kaul, 2013) $434874$ $2$ $1$
Ethylene (Fonollosa, 2015) $4178504$ $18$ $5$
Songs (Bertin-Mahieux, 2011) $515345$ $89$ $15$
Table 7. Experimental evaluation of error $\psi$ on real-world datasets, including standard deviations. Here we are in the setting where $\varepsilon=0.05$.
Dataset Local LTM Laplace LTM Gauss Central
Power $(1.206\pm 0.651)\times 10^{-2}$ $(4.036\pm 2.448)\times 10^{-3}$ $(2.108\pm 2.400)\times 10^{-6}$ $(1.763\pm 0.621)\times 10^{-6}$
Elevation $(1.193\pm 0.896)\times 10^{-1}$ $(1.520\pm 4.484)\times 10^{-2}$ $(4.718\pm 4.201)\times 10^{-8}$ $(4.465\pm 4.836)\times 10^{-8}$
Ethylene $(5.877\pm 1.256)\times 10^{-4}$ $(5.821\pm 1.074)\times 10^{-4}$ $(2.420\pm 0.486)\times 10^{-6}$ $(7.901\pm 1.909)\times 10^{-6}$
Songs $(5.617\pm 0.283)\times 10^{-7}$ $(5.629\pm 0.311)\times 10^{-7}$ $(5.556\pm 0.347)\times 10^{-7}$ $(5.613\pm 0.324)\times 10^{-7}$
In contrast to the central model of differential privacy, the performance of differentially private outputs in the local model decreases with increasing number of clients $n$. For regression and matrix factorization problems, our theoretical results for the LTM and the Gaussian mechanism suggest that this performance decrease can be avoided without sacrificing privacy under reasonable assumptions. It is natural to ask for which parametrizations of the Gaussian mechanism the performance remains acceptable. If more noise still yields good performance, we can tolerate more corrupted clients. Therefore, we evaluate empirically how the error of ridge regression and low-rank approximation develops as $n$ grows in practice and against synthetic benchmarks.
The experiments described here thus aim to demonstrate that, as the number of clients increases, the asymptotic error of differential privacy mechanisms in the LTM is between the error in the central and local models.
Setup. Our Gaussian mechanism in the LTM adds noise sampled from $\mathcal{N}(0,\sigma^{2})$ to every entry in input data matrix $\mathbf{A}$. The variance $\sigma^{2}$ depends on privacy parameters $\varepsilon$ and $\delta$ and is chosen proportional to $n^{-p}$ for $p\in[0,1]$. This enables us to interpolate between the local model ($p=0$) and the LTM ($p=1$). For each dataset we measure error for the central model, the Laplace mechanism and the Gaussian mechanism with $p\in\{1,0.9,0.8,0.7,0.6,0.5,0\}$, by reporting the average error over $20$ runs of the algorithms per dataset. The Laplace mechanism in the LTM on the other hand adds the sum of $m$ samples from $\text{Lap}(0,1/\varepsilon)$ to every entry in $\mathbf{S}\mathbf{A}$ in order to investigate the $\delta=0$ regime in the LTM.
As a baseline for low-rank approximation in the central model, we implemented MOD-SULQ (Chaudhuri et al., 2012), an approach where $\mathbf{A}^{T}\mathbf{A}$ is perturbed by adding Gaussian noise. We measure the error in terms of the excess risk;
where $\mathbf{X}_{\text{OPT}}$ denotes the optimal solution and $\mathbf{X}^{\prime}$ denotes the solution after adding noise to the training data. We vary the privacy parameter $k$ based on the dataset at hand.
For linear regression in the central model, we implemented the so-called Sufficient Statistics Perturbation (SSP) algorithm (Vu and Slavkovic, 2009). Due to its singular dependency on a multiplicative error, we use approximation factor as an error measure:
Datasets. We evaluate our mechanism on synthetic and real datasets, enabling us to vary the number of sampled points $n$ and interpolating between datasets of different sizes.
The synthetic datasets for low-rank approxiamtion have a large spectral gap from the $k$th to the $(k+1)$st singular value with random bases, which emphasizes the performance difference between the various private mechanisms. We generate synthetic data by first producing a matrix $\mathbf{A}^{\prime}$ where every entry is sampled from $\mathcal{N}(0,1)$ and then changing its singular values such that there are exactly $k$ big ones and the rest are small. More specifically, we set $\mathbf{A}=\mathbf{U}^{\prime}\mathbf{\Sigma}\mathbf{V}^{\prime}$, where $\mathbf{U}^{\prime}\mathbf{\Sigma}^{\prime}\mathbf{V}^{\prime}$ is the SVD of $\mathbf{A}$ and $\Sigma$ is a diagonal matrix with the first $k$ values set to $\sqrt{n/k}$ and the rest set to $1/n$. We vary parameter $k\in\{5,10\}$ for $d=50$.
We generate synthetic data for regression by first sampling every entry in $\mathbf{A}$ from $\mathcal{N}(0,1)$, then sampling $\mathbf{x}_{\text{OPT}}$ such that all entries are sampled from $\mathcal{N}(0,\mu^{2})$, and finally setting $\mathbf{b}=\mathbf{A}\mathbf{x}_{\text{OPT}}$. We vary parameters $d\in\{3,10,50\}$, and $\mu^{2}\in\{1,n,n^{2}\}$.
For all combinations of parameters, we generate $17$ synthetic datasets of sizes $\{1000^{i\cdot 1.5}|i=0,\dots,16\}$. For each of those we then measure error for the central mechanism, the Laplace mechanism and the Gaussian mechanism with $p\in\{1,0.9,0.8,0.7,0.6,0.5,0\}$, by running the algorithms $20$ times per dataset and reporting the average error.
In addition to synthetic datasets, we also evaluated our mechanism on $4$ datasets from the UC Irvine Machine Learning Repository. Table 6 provides their number of entries $n$ and dimensionality $d$, as well as the rank $k$ we chose for low-rank approximation experiments.
  • The first dataset consists of electric power consumption measurements in one household (Hebrail and Berard, 2012) and the feature we try to predict is sub_metering_3. We ignore the date and time features and the data points that had missing values. This leaves us with $6$ features (plus the one we are predicting) and $2049280$ data points.
  • The Elevation dataset (Kaul, 2013) consists of $434874$ open street map elevation measurements from North Jutland, Denmark. We predict the elevation from the longitude and latitude features.
  • The Ethylene dataset (Fonollosa, 2015) contains recordings of sensors exposed to a mixture of gas. We trained on the part where the sensors were exposed to a mixtures of Ethylene and CO in air. The feature we are predicting is the last one TGS2620, which leaves us with $d=18$ and $n=4178504$.
  • The Songs dataset (Bertin-Mahieux, 2011), consists of $89$ audio features that are meant to predict the release year of a song.
Low-Rank Approximation. Figure 2 shows our results for two privacy regimes ($\varepsilon\in\{0.1,0.5\}$) with synthetic data and Table 7 shows the error for $\varepsilon=0.05$ with real-world datasets. In all chosen parameter settings, we observe that as $n$ grows, the error in the LTM asymptotically approaches the error in the central model, both on real and synthetic data sets. On real-world datasets, our Gaussian approach performs significantly better than the Gaussian mechanism in the local model and is very close to the central MOD-SULQ (Chaudhuri et al., 2012) mechanism. Both with synthetic and real-world datasets, our Laplace approach performs better than the local model, though not as significantly as the Gaussian approach. However, this is the price of achieving pure $\varepsilon$-differential privacy with $\delta=0$.
@@ 6.2. Interpolation between Local and Central DP
Dataset Local Our Central
Power $2.364\pm.007$ $1.055\pm.000$ $1\pm.001$001
Elevation $1.939\pm.008$ 1.000±.000plus-or-minus1.000.0001.000$\pm$.$0$000 1.000±.000plus-or-minus1.000.0001.000$\pm$.$0$000
Ethylene $3.125\pm.002$ 1.000±.000plus-or-minus1.000.0001.000$\pm$.$0$000 1.000±.000plus-or-minus1.000.0001.000$\pm$.$0$000
Songs $20.64\pm.016$ 1.000±.000plus-or-minus1.000.0001.000$\pm$.$0$000 1.000±.000plus-or-minus1.000.0001.000$\pm$.$0$000
Ridge Regression. Figure 3 provides our experimental results based on synthetic data, and Table 8 shows the error resulting from real-world datasets. As expected, our approach performs asymptotically better than the Gaussian mechanism in the local model, but worse than SSP in the central model. We also found that as $n$ increases, the error of our approach asymptotically approaches the error in the central model. For $p=0.9,0.8,0.7,0.6$, the error eventually decreases for a sufficiently large $n$ for the privacy regimes we considered ($\varepsilon\in\{0.01,0.03,0.05,0.1\}$).
Table 8 shows the results we got by applying the local, our mechanism and SSP on real-world data. They include standard deviations, though for many of the results those are so small that they appear as 00 in the table. We found, that our mechanism performs significantly better than the Gaussian mechanism for local privacy. In most it even performs so well that there is basically no difference to the central model.
Varying $\lambda$
In Figure 4, we provide plots for the same setting of parameters as on Figure 3 ($d=10$ and $\mu^{2}=n$). Though, we also vary $\lambda\in\{1,10,100\}$ here. The experiments show that varying $\lambda$ does not change the asymptotic behavior of any of the mechanisms we investigated, in the sense that a variance proportional to $n^{-0.6}$ or lower eventually had decreasing error, while a variance proportional to $n^{0.5}$ produced increasing errors. The increase of $\lambda$ does however produce a significantly lower error in all settings (note the $y$-axes on the figure). Our theoretical results likewise predict this behaviour, as while the error bounds of Theorem 5.8 improve with increasing $\lambda$, they do not affect the dependency on $n$. Thus, we view the experiments as a confirmation that the theoretical bounds, while potentially improvable, express the correct asymptotic relationship between parameters and approximation bounds.
Acknowledgements.